A bacteria triples every 15 minutes. If there were originally 45 cultures in the lab, write the exponential function then determine how many cultures there are after 5 hours.


1
Expert's answer
2021-03-08T06:10:51-0500

For every hour, the bacteria multiple by 4t periods

The exponential function for number of cultures after t hours = 45 × 34t


Number of culture after 5 hours = 45 × 34×5 = 45 ×320

= 1.57×1011